### v1

# RADIO Amplified: Improved Baselines for Agglomerative Vision Foundation Models

Abstract
Agglomerative models have recently emerged as a powerful approach to training vision foundation models, leveraging multi-teacher distillation from existing models such as CLIP, DINO, and SAM. This strategy enables the efficient creation of robust models, combining the strengths of individual teachers while significantly reducing computational and resource demands. In this paper, we thoroughly analyze state-of-the-art agglomerative models, identifying critical challenges including resolution mode shifts, teacher imbalance, idiosyncratic teacher artifacts, and an excessive number of output tokens. To address these issues, we propose several novel solutions: multi-resolution training, mosaic augmentation, and improved balancing of teacher loss functions. Specifically, in the context of Vision Language Models, we introduce a token compression technique to maintain high-resolution information within a fixed token count. We release our top-performing models, available in multiple scales (-B, -L, -H, and -g), alongside inference code and pretrained weights.
Links:Code (on GitHub) | Models (on Hugging Face)
11footnotetext: Equal Contribution
1 Introduction
Refer to captionRefer to caption
Figure 1: Overview of the main contributions: From left to right, we introduce a multi-resolution training framework that enables our student model to maintain accuracy across all resolutions; we propose using token merging to better retain fine-grained details while merging uninformative tokens; we improve upon the AM-RADIO baseline across all benchmarks, with significant gains on VLM tasks.
The rise of specialized vision foundation models has created a need for methods to consolidate knowledge from multiple models into a single one. SAM-CLIP [39] addresses this challenge by combining SAM [19] and CLIP [27] to integrate capabilities from both. In another approach, AM-RADIO [30] introduces label-free knowledge distillation from multiple teacher models, enabling knowledge transfer without direct supervision. UNIC [32] adds intermediate teacher matching projectors and dynamic teacher selection. Theia [33] aims to facilitate robot learning by distilling insights from multiple vision teachers. Meanwhile, PHI-S [29] studies the importance of normalizing the distinct teacher distributions to simplify their balancing. Relatedly, Eagle [34] leverages a mixture of vision encoders to achieve inference-time knowledge aggregation within the context of vision-language models (VLMs).
Despite these advancements, this growing body of work on knowledge agglomeration still leaves open several critical challenges:
  • Resolution balancing: Teacher models operate at varying resolutions due to different architectures and training goals, creating feature granularity inconsistencies. Effective techniques are needed to balance these resolutions in the student model to capture both fine details and broader abstractions.
  • Teacher distribution balancing: Existing models have different distribution moments and the distillation process should account for this to prevent biased learning.
  • Generating multi-resolution features for diverse applications: Vision models support various applications requiring different feature resolutions, from image captioning to dense segmentation. A VFM that can flexibly produce features at any resolution could serve multiple tasks, reducing the need for separate models, and opening up new opportunities.
In our analysis of existing agglomerative models, we study and propose a fix for the notable “mode switch” phenomenon in AM-RADIO, where feature distributions shift significantly based on input resolution (Section 3.1). Specifically, low-resolution inputs yield DINO-like features, while high-resolution inputs produce SAM-like features. We trace this behavior to the student learning different teachers at different resolutions during training. In section 4.2, we introduce a solution to stabilize these mode switches, achieving strong resolution robustness and improved quality scaling.
Armed now with a vision encoder that works best at high resolution, we next look toward integrating it into VLMs. In the case of downstream applications with vision-language models, a common pitfall is the number of output tokens/features. By processing images in high resolution, most methods will return more image tokens and will result in quadratic complexity of attention in VLMs. High resolution processing is important, and in order to preserve this information, we propose to apply ToMeSD [3] in section 4.7 to decouple the vision encoder resolution from the number of patches used by the VLM.
Our main contributions are as follows:
  • A multi-resolution training strategy that fixes mode switching and allows for fully flexible input resolution.
  • A comprehensive study of feature selection for a range of downstream tasks.
  • A new method to compress visual features, enabling integration with language models while preserving essential information. We demonstrate that our proposed vision encoder disproportionately benefits from this.
Model Goal ImageNet1K Segmentation Probe3D[11] Vision-Language (VILA[20]) SAM [19]
Zero-shot k-NN ADE20k Depth SurfNormals TextVQA ChartQA DocVQA InfoVQA OCRBench COCO
AM-RADIO-L* Baseline 78.35 84.01 50.03 82.63 59.94 57.01 15.6 24.19 28.88 25.60 73.77
+ multi-res Eliminate modes 81.21 84.09 52.84 82.39 61.04 59.44 16.6 33.99 29.02 29.40 75.49
- OpenAICLIP + SigLIP Better VLM 81.01 84.68 52.95 84.7 60.1 64.42 25.28 45.80 30.60 36.00 74.19
+ ViT-H Bigger backbone 82.51 85.81 53.97 85.7 62.5 65.88 25.96 49.74 35.17 40.90 76.14
+ Token Merging Improve VLM - - - - - 69.74 30.40 52.33 36.24 42.90 -
Table 1: Ablation Results. For Probe3D, Depth and Surface Normals metrics are averaged over the buckets defined in the paper. For SAM COCO we use the “instance all” bucket. Each incremental change we introduce leads to improved metrics. *Our reproduction.
2 Background
2.1 Baseline Model
Following AM-RADIO[30], our model consists of a Vision Transformer (ViT)[10] backbone, including CPE for multi-resolution support. We use DFN CLIP [13], OpenAI CLIP [27], DINOv2-g-reg [9] and SAM-H [19] as teachers. For each teacher to distill from, we augment the backbone with an adaptor for the summarization token, and an adaptor for the patch tokens. Our adaptor is a 2-layer MLP with a LayerNorm[1] and a GeLU[16] in between. We train the student using teacher features generated by inferring images of the DataComp1B[14] dataset. We train the model with a batch size of 1024+128 for 600k iterations. Training is split into two concurrent partitions: one partition trains the student at a resolution of $432^{2}$ against CLIP and DINOv2 teachers with batch size 1024, while the other partition trains the student at a resolution of $1024^{2}$ against SAM with a batch size of 128.
2.2 Evaluation Framework
Evaluating the quality of a foundation model is a daunting task due to the wide spectrum of downstream applications. We employ a rigorous evaluation framework to identify areas for improvement and guide design decisions:
  • Image-level reasoning: We focus on ImageNet-1k [31] Top-1 classification accuracy using (i) zero-shot [27] classification using a pretrained and frozen language model; and (ii) k-NN [4] classification.
  • Pixel-level semantic segmentation: We linearly probe the frozen backbone features for semantic segmentation.
  • Instance segmentation: We evaluate our model on the COCO [18] dataset.
  • 3D understanding: We consider depth estimation, surface normals estimation, multi-view correspondence, and semantic correspondence, following [11].
  • Vision-Language reasoning: We pair our backbone with an LLM to evaluate Vision-Language modeling.
3 Challenges
3.1 Achieving Multi-Resolution Robustness
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Zero-Shot Accuracy and teacher matching fidelity metrics as a function of input resolution. AM-RADIO, and our “Multi-Stage Stage 3” model exhibit mode switching behavior. “Multi-Stage Stage 2” doesn’t exhibit the behavior, owing to all teachers being trained solely at low resolution, however quality at high resolution quickly degrades. “Multi-Resolution” training fixes the mode switch, and allows the model to be strong across a larger range of resolutions compared to Stage 2.
Although AM-RADIO [30] supports a wide range of input resolutions, we noticed poor performance on high-resolution benchmarks. As illustrated in Figure 2, Zero-Shot Top-1 accuracy drops sharply after inputs are larger than roughly 512px. This showcases an undesirable property of this model, which is that it doesn’t reliably scale to high-resolution images. AM-RADIO hints at this issue in its conclusion: the model exhibits a “mode switching” behavior as resolution increases. This prompted us to visualize the features and investigate the effect of a resolution increase. Figure 3 illustrates the issue well: at resolutions lower than or equal to $512^{2}$, the features most closely resemble those of DINOv2 and appear to be expressive of depth and semantic content. At higher resolutions, the model starts to behave more like SAM, with features that show sharp delineations around contours, with homogeneous contents within. This behavior is intuitive given the fact that in the high-resolution regime the student only sees SAM features. Table 2 shows a quantified measure of feature variance across scales (see implementation in appendix). DINOv2 exhibits much smaller scale variance.
Figure 3: Visualizations of patch token features for different input resolutions. We use PCA to project patch tokens into a 3D-space representing RGB colors. Top: From left to right: input image, visualizations of DINOv2, AM-RADIO at $256^{2}$, $384^{2}$, $512^{2}$, $768^{2}$ and $1024^{2}$ resolutions, and SAM. Bottom: RADIO-AMP (ours) features. As resolution increases, AM-RADIO features look increasingly similar to those of SAM, even exhibiting VitDet windowing artifacts. RADIO-AMP features are much more self-consistent and less noisy. More samples can be found in the Appendix.
3.2 Token Count
Dense downstream tasks, such as pixel-level semantic segmentation and depth estimation, benefit from receiving fine features from a vision encoder. However, for vision-language models like LLaVA [21], where visual tokens are integrated into a sequence of text embeddings, an excessive number of vision tokens can negatively impact performance or lead to sequence overflows. Pixel unshuffling, as proposed in InternVL1.5 [6], addresses this by reducing the number of vision tokens, grouping $2\times 2$ tokens along the channel dimension. However, this approach cannot adapt to varying information densities within an image. For instance, a text document might have large areas of white background that should be more strongly compressed than areas containing text.
4 Method
4.1 Scale Equivariance
4.1.1 A Measure of Scale Equivariance
We define a measure of scale equivariance for a set of multi-scale features by normalizing and interpolating all features to the scale of the lowest resolution, then calculating the spatial average of the variance along the scale dimension. Our formula can be found in appendix A.5.
As can be seen in Table 2, our baseline student exhibits much worse scale equivariance than that of DINOv2 (the only teacher that is capable of operating at multiple resolutions, hence our only point of comparison).
4.1.2 Tiling
Tiling is commonly employed in VLMs (LLaVa-NeXT[21], InternVL1.5[6]) as a way to enable high-resolution inference when the vision encoder only supports fixed-resolution inputs. Thus it could be argued that scale equivariance is of little importance, given that it is possible to resort to tiling instead of increasing input image resolution. However tiling incurs additional challenges:
  • Tiling makes resolution increases very coarse.
  • As the number of tiles increases, the vision encoder sees an increasingly small subset of the full image, limiting its ability to reason about the full context, or even know what it’s looking at entirely.
Figure 4: Qualitative comparison of tiling and high-resolution inference. Tiling emulates high-res SigLIP[44] inference. Top: From left to right: input, SigLIP features using 1, $2\times 2$ and $4\times 4$ and $5\times 5$ tiles. Bottom: Single-pass inference using RADIO-AMP (ours) at equivalent image sizes. SigLIP features vary significantly across different tiling arrangements.
Figure 5: Visualizations of SigLIP and RADIO-AMP model features before and after ToMe compression. From left to right: input image; PCA visualization of features; merged tokens; PCA visualization of unmerged tokens. RADIO-AMP is better able to compress background regions, leaving more tokens available for semantically rich features.
In Figure 5 we use SigLIP [44], a popular choice of Vision Encoder (Cambrian [38]), and show its features for multiple tiling arrangements. We notice qualitative inconsistency in the way features get represented at multiple scales. We can also apply our measure of scale equivariance, albeit rather coarsely, to quantify this inconsistency (Table 2). In both cases we notice the detrimental effect of tiling on scale equivariance. Tiling also incurs a substantial increase in the number of vision tokens, by a factor of up to 12×12$\times$12 × ([6]), with immediate effects on VLM latency.
Finding 1. High-resolution inference through tiling causes the vision encoder to lose global context and exhibit poor scaling equivariance.
4.2 Multi-Resolution Training
The previous observations motivate a change in the training schedule: with multi-resolution training we enable the student to learn from all teachers across multiple resolutions. This is easily achieved with DINOv2, since DINOv2 can infer images at any resolution. For CLIP teachers, we feed them with images at the teacher’s native resolution, and we feed the student with images at multiple resolutions. We interpolate student features down to the resolution of the teacher’s features before applying the loss function. SAM presents an extra challenge, as noted in AM-RADIO[30], since interpolating SAM features significantly deteriorates their quality. Therefore in order to train our student against SAM at lower resolutions than $1024^{2}$, we pad smaller images to $1024^{2}$, then crop SAM features to the effective size of the unpadded image.
Training the student for 600k iterations with a multi-resolution setup is costly. Thus we break down training into three stages:
  • In a first stage, we train the student at low resolution ($256^{2}$) for 300k iterations.
  • In a second stage, we train the student at medium resolution ($432^{2}$) for 300k iterations.
  • In the third stage, we train the student simultaneously at $432^{2}$ and $1024^{2}$ resolutions for 300k iterations.
Finding 2. For the student model to be consistently accurate across resolutions, it is sufficient to match all teachers at all resolutions, and to train at two resolutions simultaneously in the final training stage.
Equivariance $\downarrow$
Model Fine Scale Coarse Scale
DINOv2-g-reg 0.126 0.178
AM-RADIO-L 0.357 0.476
RADIO-AMP-B 0.102 0.168
RADIO-AMP-L 0.102 0.165
RADIO-AMP-H 0.119 0.193
RADIO-AMP-L (tiling) N/A 0.369
SigLIP (tiling) N/A 0.623
Table 2: Scale equivariance, computed over a small dataset of natural images (shown in Appendix). Fine: we scale image resolutions from $384^{2}$ to $1568^{2}$ by increments of the model’s patch size. Coarse: we scale images from $384^{2}$ to $1920^{2}$ by multiples of $384^{2}$. For the models that are denoted as “tiling”, we perform inference over tiles of $384^{2}$) and re-assemble features to produce the high-resolution features.
Figure 6: Illustration of the multi-resolution training setup. The student learns from all teachers at all resolutions. High-res student features are downsampled to match the resolution of low-res teachers. High-res teachers are provided with mosaics of smaller images and their output features are cropped out in order to train a low-res student.
4.3 Mosaic Augmentation
The training schedule described in Section 4.2 involves running SAM inference on padded images, using cropped features to train the student against SAM at low resolution. This approach incurs a substantial computational cost, as SAM is the most resource-intensive teacher.
To improve efficiency when training a student at a resolution $\leq 512^{2}$ against SAM, we can instead create a mosaic of $k\times k$ images, with $k=\left\lfloor\frac{1024}{x}\right\rfloor$ and $x$ being the student resolution, resulting in a single $1024^{2}$ image. We then perform SAM inference on this mosaic and extract $k^{2}$ individual feature maps to train the student. Mosaic augmentation includes padding as needed to maximize efficiency. For example, to train a student at $432^{2}$ resolution, we can create a $2\times 2$ mosaic with 80-pixel padding around each image. Figure 7 shows sample mosaic augmentations under 256 and 432 student resolutions. Qualitatively, we observe cleaner features after applying mosaic augmentation, which we believe is due to the increased diversity in image positions, helping to reduce positional encoding artifacts. We also show SAM’s PCA features for these mosaic images in figures A5 and A4 in the appendix.
Figure 7: Sample mosaic augmentations. Grid lines represent $16^{2}$ patches (only shown here for visualization purposes). Left: A $4\times 4$ arrangement with a $1024^{2}$ image comprising $16\times 256^{2}$ sub-images. Right: A $2\times 2$ arrangement comprising $4\times 432^{2}$ sub-images, with 80 pixels of padding around each.
Finding 3. Mosaic augmentation greatly reduces the training cost associated with learning from high-resolution teachers and eliminates the need for feature interpolation. Student quality is even improved with this optimization.
4.4 Teacher Loss Balancing
PHI-S [29] highlights the significant variations in activation magnitudes among vision foundation models, observing that SAM’s activations tend to overshadow those of CLIP and DINOv2 models. We adopt the PCA-Hadamard Isotropic Standardization (PHI-S) method, achieving improved balance among teacher losses. PHI-S rotates teacher activations to evenly distribute variance across all channels and then scales them to obtain unit variance. This process can be easily reversed by projecting the student activations back into each teacher’s original feature space. PHI-S enhances training stability and overall benchmark performance. For a given teacher feature map $\mathbf{X}$ with embedding size $C$, PHI-S applies the following transformation:
$\mathbf{X}_{i}^{\prime}=\phi_{i}^{-1}\mathbf{R}_{i}\mathbf{X}_{i},\quad\mathbf{R}_{i}=\mathbf{H}_{C}\mathbf{U}_{i}^{\intercal},\quad\phi_{i}=\sqrt{\frac{1}{C}\sum_{j}^{C}\lambda_{j}}$ (1)
Where $H_{C}$ is a normalized Hadamard matrix of dimension $C$, $\lambda_{j}$ are the Eigen values of the covariance matrix $\mathbf{\Sigma}\left[\mathbf{\mathbf{X}}\right]$, and $\mathbf{U}$ are the corresponding eigenvectors. $\phi_{i}$ and $\mathbf{R}_{i}$ are specific to the $i$th teacher.
As a starting point, we define a measure of fidelity, similar to that in the classification-distillation literature ([37, 2]), however we do so without the use of labels or explicitly produced distributions over classes. Instead, since our loss objective is to directly match the features of the teachers, we have
[…]
with $f(X)$ being the student feature distribution, and $t_{i}(X)$ being the $i$th teacher distribution. This function represents the ratio of the target distribution variance to the student’s estimation error variance. A value of $\leq 1$ means random sampling from the teacher distribution would be better, and $\infty$ would be perfect matching. We show the results of this in table 3, where it is apparent that the baseline allocates too much energy to matching SAM due it its disproportionately large distribution variance, consistent with [29]. The errors relative to the variance are overall smaller when applying PHI-S.
Teacher $\phi^{2}$ MSE F[X]
Baseline PHI-S Baseline PHI-S
DFN CLIP 5.831E-4 5.100E-4 2.418E-4 1.143 2.411
OpenAI CLIP 0.820 0.570 0.525 1.438 1.563
DINOv2-g 1.729 0.222 0.206 7.799 8.377
SAM 27.263 3.719 5.313 7.331 5.132
Geometric Mean Baseline PHI-S
3.114 3.568
Table 3: Fidelity results for each teacher, comparing between the baseline and PHI-S. Higher values are better.
Finding 4. PHI Standardization helps balance the energy spent learning from each teacher.
4.5 Is SAM a good teacher?
SAM [19] has been a controversial choice in the recent agglomerative models literature. AM-RADIO [30] struggled to prove that its inclusion improved any metrics. Theia [33] specifically ablated whether to include SAM, and found that it degraded their metrics. UNIC [32] opted as well not to include SAM. Based on the findings with PHI-S [29], and our confirmation of imbalance in section 4.4, it seems that a major problem with SAM may just have been that interpolating its features is a really bad thing, and also that the distribution was extremely unbalanced; something that PHI-S corrects. We chose to re-run the study of whether SAM is a good teacher now that we have a new bag of tricks. In particular, we run the first two stages of multi-stage training, we use the mosaic augmentation for SAM in both of these stages, and we apply PHI-S to all teachers. We show the results in table 4 where it is clear that including SAM has negligible (but positive) impact on our classification benchmarks, and strong positive effects on dense tasks such as semantic segmentation and 3D probing. SAM’s inclusion also enables novel opportunities such as those found in VideoSAM [15] where they employ AM-RADIO’s SAM adaptor and backbone simultaneously for video segmentation.
Variant Zero Shot kNN ADE20k Depth SNorm MultiView SPairs
Stage 1
No SAM 79.38 83.17 50.27 82.54 61.03 58.12 51.97
SAM 79.37 83.29 51.14 82.60 61.88 58.91 52.49
Stage 2
No SAM 80.43 83.83 50.24 83.29 61.43 58.98 54.72
SAM 80.47 83.92 51.36 83.17 62.80 61.36 54.66
Table 4: Ablation on whether to include SAM in the teacher set for the first two stages of multi-stage training.
Finding 5. All teachers are beneficial, including SAM, despite recent trends. It also has broad downstream applicability, granting our student the same abilities.
4.6 Partitioning
In PHI-S [29], the authors opted to put teachers in their own partitions, which reduces the teacher overhead (as the per-teacher batch size is reduced). However, the paper does not ablate whether this choice came with model quality consequences. In Table 5 we study the number of partitions for the first two stages of training. We find that fewer partitions is strongly better for summarization tasks, and less clear for dense tasks.
# Zero Shot kNN ADE20k Depth SNorm MultiView SPairs SAM
Stage 1
1 79.37 83.29 51.14 82.60 61.88 58.91 52.49 71.51
2 78.70 83.00 50.88 82.32 60.78 58.49 52.32 71.29
4 77.93 82.61 51.09 82.89 61.26 58.47 52.54 71.02
Stage 2
1 80.47 83.92 51.36 83.17 62.80 61.36 54.66 73.62
4 78.82 83.41 51.34 82.99 61.58 59.80 55.96 73.03
Table 5: Ablation on number of partitions for the first two stages of multi-stage training.
Finding 6. Minimizing the number of partitions seems to be beneficial, assuming you can afford the teacher overhead. Under compute constraints, partitioning is an effective strategy to reduce the overhead.
4.7 Token Compression
Figure 8: Illustration of bipartite matching [3] using $2\times 2$ strided pattern with $r=9$. A: The original tokens, with the red ‘T#’ squares being assigned as targets, and the green ‘S#’ squares being assigned as sources. B: We compute the affinity between each source and target. We only consider the maximum affinity for each source (shown in yellow). We find the $r$ highest affinity yellow squares, and merge those into their respective targets. C: The output tokens with merged values when a given ‘T#’ was assigned one or more sources. D: The final $7$ tokens are fed to the LLM. Reconstructed Viz: From (C), we can visualize the compressed original feature map by broadcasting the merged tokens to all of the source locations.
Inspired by ToMeSD[3], we evaluate the use of bipartite soft matching to merge similar tokens. We apply strided partitioning to ensure that each image region retains some representation in the compressed features. For evaluation, we track merged token indices, enabling us to unmerge tokens and measure the reconstruction error and inform hyperparameter selection. The method is illustrated in Figure 8. While ToMe recommends using attention keys for matching, we found that using the features themselves improves reconstruction error. Table A9 of the appendix summarizes our findings.
We also qualitatively evaluate token merging by visualizing feature and token merging outcomes in Figure 5. Token merging is assessed within the context of VLMs, with results reported in Table 1.
We further note that token merging can be applied to other vision encoders and demonstrate its application on SigLIP. Interestingly, token merging yields a greater relative improvement on RADIO-AMP than on SigLIP, a finding that aligns with the feature visualizations in Figure 5 and the measured reconstruction error in Table A9. We hypothesize that SigLIP’s higher feature noise makes clustering more error prone.
Res Method
Tokens/im
TextVQA
ChartQA
DocVQA
InfoVQA
OCRBench
512 $2^{2}$ Unshuffle 256 65.0 23.7 45.7 32.9 40.2
ToMe r=768 256 65.7 25.8 49.7 34.9 41.2
768 $3^{3}$ Unshuffle 256 65.8 25.9 49.7 35.2 40.9
ToMe r=2048 256 69.7 30.4 52.3 36.2 42.9
1024 $4^{4}$ Unshuffle 256 66.1 23.8 46.6 34.1 38.9
ToMe r=3840 256 69.3 28.1 49.1 34.6 42.7
ToMe r=3584 512 68.9 31.3 53.9 37.0 44.6
Table 6: VLM benchmarks across various compression methods. All configurations produce 256 vision tokens per image, except for the last row (512 tokens). Token Merging outperforms Pixel Unshuffling on all benchmarks.
Finding 7. Token Merging is very effective at retaining the most diverse information under high compression ratios.
4.8 Feature Selection
For each image, our foundation model outputs a summary vector along with patch tokens at a granularity of one per $16^{2}$ input pixel block. For image-level tasks such as classification, search, or curation, the summary vector provides a rich embedding. For dense tasks, such as segmentation or 3D understanding, the patch tokens are a natural choice. However, as demonstrated in previous work, incorporating additional intermediate activations further enhances performance. For example, [11] uses a Dense Prediction Transformer (DPT) head [28] for 3D reasoning, while [43] averages multiple ranges of intermediate activations to feed into an LLM for VLMs.
In this section, we investigate various feature selection methods and present the results in Table 7. We experiment with sparse feature selection (selecting activations from individual layers throughout the model) and dense feature selection (aggregating information across all layers by averaging groups of layer activations). We examine the impact of feature selection in conjunction with different downstream heads (linear or DPT probe).
Our findings show that a linear probe alone is insufficient to leverage additional information from intermediate layer activations. However, when a DPT head is employed, it effectively incorporates this additional information. We note, however, that it is challenging to disentangle the benefits provided by additional feature information, and those of the extra learnable parameters of the DPT head. Unlike [43], we do not observe a positive impact of using dense features in VLMs.
Finding 8. Intermediate layer activations greatly benefit downstream tasks if a non-linear transformation is employed.
Layers Aggre- Head ADE20k Depth Surf Overall
gation Normals
31 N/A Linear 52.47 82.9 57.0 61.215
7-15-23-31 Sparse Linear 52.99 82.5 59.6 62.03
(0-9)-(10-19)-(20-30)-31 Dense Linear 52.96 82.7 59.5 62.03
15-31 Sparse Linear 52.90 83.1 59.6 62.12
(0-15)-(16-30)-31 Dense DPT 54.27 85.4 60.7 63.65
15-31 Sparse DPT 54.58 84.6 61.0 63.70
7-15-23-31 Sparse DPT 55.19 85.9 61.6 64.46
(0-9)-(10-19)-(20-30)-31 Dense DPT 54.28 85.5 62.3 64.08
3-7-11-15-19-23-27-31 Sparse DPT 54.42 86.7 62.8 64.58
TextVQA ChartQA DocVQA InfoVQA OCRBench
Last N/A 63.6 23.4 47.0 33.8 42.0
7-15-23-31 Sparse 63.2 24.1 47.2 34.3 40.3
(0-9)-(10-19)-(20-30)-31 Dense 63.5 23.1 47.0 33.5 40.2
Table 7: Study on the effect of RADIO-AMP-H feature selection. For “dense” aggregation, the numbers in brackets indicate the range of layers from which the average is calculated. Top: Semantic segmentation, depth estimation, surface normals estimation. Bottom: VILA benchmarks. Pixel-level tasks exhibit a clear preference for more layers and non-linear heads, while VLM tasks seem mostly neutral to this choice.
4.9 SigLIP Teacher
Building on previous related work ([38], [20]), we replace OpenAI-CLIP [27] with SigLIP [44] and observe significant improvements in VLM tasks, as shown in Table 1.
5 Experiments
Model Params ADE20k NYUd Pascal Context
SemSeg SemSeg Depth Normal SemSeg Parsing Saliency Normal
[…]
SAM-CLIP[39] 86M 38.4
Theia [33] 86M 35.55 38.90 0.6377 24.11 69.84 60.67 80.63 16.94
UNIC-B[32] 86M 39.6 42.21 0.6172 22.78 75.90 62.85 81.84 15.78
RADIO-AMP-B (ours) 98M 48.94 57.19 0.4980 20.04 81.75 71.49 81.26 16.10
UNIC-L[32] 307M 48.30 58.56 0.4916 19.34 81.82 72.24 79.21 17.35
RADIO-AMP-L (ours) 320M 52.95 61.42 0.4577 18.57 82.87 74.32 81.65 16.15
UNIT[46] 632M 50.19 61.44 0.4538 19.07 82.36 73.34 78.70 17.98
AM-RADIO-H[30] 653M 51.36 62.76 0.4339 18.43 82.78 74.42 78.48 17.53
RADIO-AMP-H (ours) 653M 53.97 63.82 0.4353 17.67 83.43 75.75 81.19 16.16
UNIT[46] 1B 50.63 61.62 0.4527 19.07 82.20 73.22 78.51 17.96
RADIO-AMP-g (ours) 1.1B 54.56 63.46 0.4318 17.08 82.65 75.18 80.08 17.06
Table 8: Comparison of dense task benchmarks against other agglomerative models. † Results unavailable due to lack of access to SAM-CLIP model weights.
Vision Encoder
Resolution
Tokens/im
TextVQA
ChartQA
DocVQA
InfoVQA
OCRBench
GQA
POPE
MME
SEED(I)
MMMU
AI2D
Average
OpenAI-CLIP [27] $336^{2}$ 196 63.2 49.2 43.4 27.4 431 62.4 85.7 1517.9 72.7 43.4 77.6 58.55
AM-RADIO-H [30] $512^{2}$ 196 63.6 51.7 40.1 26.2 392 63.4 87.3 1570.6 73.6 43.9 78.9 58.76
AM-RADIO-H [30] $768^{2}$ 196 62.8 49.9 36.7 24.8 365 62.7 87.0 1517.2 71.7 41.9 73.4 56.66
SigLIP-SO400M [44] $384^{2}$ 196 64.8 55.7 47.9 29.2 452 63.1 85.4 1632.9 72.9 41.1 80.7 60.95
RADIO-AMP-L (ours) $768^{2}$ 196 66.7 56.4 49.2 29.8 441 63.4 87.6 1592.4 74.0 43.3 79.2 61.21
RADIO-AMP-H (ours) $768^{2}$ 196 68.8 61.6 54.2 30.8 498 63.8 87.3 1562.3 74.1 44.1 82.5 63.19
RADIO-AMP-g (ours) $768^{2}$ 196 69.6 57.0 53.3 30.9 479 63.4 87.5 1606.4 74.3 41.3 80.8 62.39
Table 9: Vision Encoder Comparison in VILA[20]. We used MN-Minitron-8B[36] as LLM and the LLaVA1.6[21] data mixture. We use Token Merging to compress all visual features down to 196 tokens. From left to right we report: image resolution, number of tokens per image, TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, MMMU (validation) accuracy, AI2D (no mask) accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
5.1 Ablation Studies
Ablation studies are summarized in Table 1. Our initial baseline is a reproduction of AM-RADIO, using a ViT-L backbone and the DataComp1B[14] dataset. Our hyperparameters are detailed in the appendix. We experiment with several variations of the training schedule, as described in Section 4.
For Image classification, we report the ImageNet-1k[31] Top-1 classification accuracy using two methods. i) zero-shot[27] classification using the DFN CLIP language model; ii) k-NN[4] classification using the embeddings of the shared backbone.
For Instance Segmentation, we use the EfficientViT[22] implementation of a COCO[18] instance segmenter to measure how well our SAM adaptor head can replace SAM.
For Semantic Segmentation, we use the MMSeg[8] framework, and train a linear probe on top of the frozen RADIO-AMP features. We use the ADE20k[45] dataset. We train the probe at three resolutions ($512^{2}$, $768^{2}$, and $1024^{2}$) and report the best result (full results in appendix).
For 3D understanding experiments we use the Probe3D[11] code base and apply a DPT head on top of the frozen backbone features. We train and evaluate on the Navi[17] dataset. We resize images to $512^{2}$ during both training and evaluation.
For Vision-Language Modeling we follow the same framework as in VILA [20]. We pair our vision encoder with a MN-Minitron-8B[36] LLM. We train all alignment/pretraining/SFT and unfreeze the vision encoder during SFT. We train the model using ShareGPT4v[5] and VFLAN[41] datasets. We leverage insights from Cambrian[38] to select the benchmarks that are most sensitive to the quality of visual features report scores on TextVQA[35], ChartQA[24], DocVQA[26], InfoVQA[25] and OCRBench[23].
5.2 Comparative Results
5.2.1 Dense Task Evaluation
[…]
Results are reported in Table 8. RADIO-AMP models exhibit a consistent improvement over previous baselines.
5.2.2 VLM Evaluation
We compare our models against state-of-the-art vision encoders: OpenAI-CLIP [27], AM-RADIO-H [30] and SigLIP-SO400M [44]. We pair the vision encoders with MN-Minitron-8B [36] and train them in VILA [20] using the LLaVA1.6 [21] data mixture. For all vision encoders, we use Token Merging to compress the visual features to 196 tokens in order to ensure a fixed cost to the Language Model. We report results in Table 9. We observe that RADIO-AMP models consistently exceed the accuracy of other vision encoders.
6 Conclusion
In this paper, we presented RADIO Amplified, a robust framework for multi-teacher vision model distillation that accommodates variations in teacher resolutions, activation magnitudes, and noise patterns. The key differentiating factor in our work is our focus on preserving accuracy across a broad range of resolutions. Our multi-resolution training approach mitigates resolution-dependent performance degradation (mode switching). Our findings show that mosaic augmentation and PHI-S effectively balance computational load and loss contributions from each teacher.
Token compression enables efficient integration with VLMs by preserving critical visual information, even at high compression ratios. Our feature selection study highlighted the advantages of using intermediate activations for dense tasks, particularly when non-linear transformations are used.
Appendix A.1 VLM Benchmarks
A.1.1 More Vision Encoder Comparisons
Vision Encoder
Resolution
Tokens/im
TextVQA
ChartQA
DocVQA
InfoVQA
OCRBench
GQA
POPE
MME
SEED(I)
AI2D
Average
OpenAI-CLIP [27] $336^{2}$ 144 63.8 27.5 48.8 33.0 414 63.0 86.7 1646.9 66.7 67.1 58.03
AM-RADIO-H [30] $512^{2}$ 256 55.9 15.7 35.2 30.6 316 60.2 85.3 1516.5 71.8 63.1 52.52
SigLIP-SO400M [44] $384^{2}$ 196 67.6 33.0 57.1 36.0 458 63.0 85.7 1605.2 74.2 68.6 61.13
RADIO-AMP-H (ours) $768^{2}$ 196 70.8 33.9 59.5 37.0 482 63.9 86.9 1613.5 75.1 66.7 62.27
RADIO-AMP-H (ours) $768^{2}$ 512 70.2 37.3 64.2 37.6 523 64.5 87.3 1587.6 75.4 67.5 63.57
Table A1: VILA benchmark results for various vision encoders. We used Qwen2-7B-Instruct as LLM and ShareGPT4v[5] and VFLAN[41] data. From left to right we report: image resolution, numbers of tokens per image, TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, AI2D accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
In Table A1, we report benchmark results for OpenAI-CLIP-336, AM-RADIO-H, SigLIP-400m, and RADIO-AMP-H. The same Qwen2-7B-Instruct LLM, training data, and hyperparameters are used across all configurations. RADIO-AMP-H is utilized in conjunction with Token Merging and configured to produce either 196 tokens per image (matching SigLIP) or 512 tokens per image (to demonstrate scaling capabilities). The results show a significant improvement when transitioning to RADIO-AMP-H, even with the same token count as the SigLIP baseline. Furthermore, additional benefits are observed when scaling up the number of vision tokens.
A.1.2 Scaling up to More Data
Vision Encoder TextVQA ChartQA DocVQA InfoVQA OCRBench GQA POPE MME SEED(I) AI2D Average
SigLIP-SO400M [44] 69.7 67.2 63.7 40.9 588 62.4 86.9 1648.5 72.0 75.0 67.90
RADIO-AMP-H (ours) 71.5 71.9 73.6 45.6 667 62.7 87.5 1664.1 77.39 74.8 71.49
Table A2: VILA benchmark results for SigLIP and RADIO-AMP-H. We used Qwen2-7B-Instruct [40] as LLM, with an improved data mixture of public datasets. From left to right we report: TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, AI2D accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
In Table A2, we report benchmark results obtained using an improved SFT data mixture, which includes data from ShareGPT4v, LLaVA Instruct, Cambrian, VFLAN, and the training sets of some benchmarks, for a total of 9.8M samples. RADIO-AMP-H is used in conjunction with Token Merging and is configured to produce 512 tokens per image. RADIO-AMP-H outperforms the SigLIP baseline on all benchmarks, except for AI2D, where it achieves a tie with SigLIP.
A.1.3 Effect of the Compression Method
Vision Encoder Compression
Tokens/im
TextVQA
ChartQA
DocVQA
InfoVQA
OCRBench
GQA
POPE
MME
SEED(I)
AI2D
Average
SigLIP SO400M [44] $2\times 2$ Unshuffle 196 63.4 23.1 41.4 30.1 339 63.8 85.4 1518.4 70.9 63.6 55.15
SigLIP SO400M [44] ToMe r=533 196 62.9 21.2 41.8 30.1 326 64.2 85.8 1537.5 71.7 63.7 55.09
RADIO-AMP-H (ours) $2\times 2$ Unshuffle 576 66.4 25.0 53.3 32.5 402 64.8 86.5 1434.0 73.4 63.9 57.77
RADIO-AMP-H (ours) ToMe r=2048 256 69.7 30.4 52.3 36.2 429 63.8 86.8 1572.4 74.6 65.1 60.04
RADIO-AMP-H (ours) ToMe r=3584 512 68.9 31.3 53.9 37.0 446 63.0 87.6 1537.7 73.9 66.0 60.31
Table A3: VILA benchmark results for various vision encoders and compression methods. We used MN-Minitron-8B as LLM and ShareGPT4v[5] and VFLAN[41] data. From left to right we report: number of vision tokens per image, TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, AI2D accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
In Table A3, we report benchmark results for SigLIP-400m and RADIO-AMP-H using different token compression methods. The same MN-Minitron-8B LLM, training data, and hyperparameters are used across all configurations. The results show no improvement when applying Token Merging to SigLIP. RADIO-AMP-H, on the other hand, demonstrates significant improvement with Token Merging, and increasing the token count from 256 to 512 provides a modest additional gain.
A.1.4 High-Resolution Inference using Tiling
Vision Encoder Resolution Compression
Tokens/im
TextVQA
ChartQA
DocVQA
InfoVQA
OCRBench
GQA
POPE
MME
SEED(I)
AI2D
Average
SigLIP SO400M [44] + Tiling Up to $13\times 384^{2}$ $2\times 2$ Unshuffle $\sim 1928$ 66.5 64.6 74.4 40.1 521 64.5 88.0 1536.1 74.4 79.3 68.07
RADIO-AMP-H (ours) $768^{2}$ ToMe r=2108 196 68.8 61.4 61.9 36.7 498 63.8 88.4 1562.6 74.1 71.9 65.49
RADIO-AMP-H (ours) + Tiling Up to $7\times 768^{2}$ ToMe r=2108 $\sim 1233$ 73.2 63.5 73.1 46.5 545 64.5 87.9 1611.6 75.3 82.4 70.15
Table A4: VILA benchmark results using tiling to emulate high-resolution inference. We used MN-Minitron-8B as LLM and the data mixture from LLaVA1.6. From left to right we report: number of vision tokens per image (calculated across all benchmarks), TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, AI2D (No Mask) accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
In Table A4, we emulate high-resolution inference using tiling [7]. While we acknowledge that there is no strict equivalence in the number of pixels (SigLIP processes an average of 1.8M pixels per sample, whereas RADIO-AMP-H processes an average of 3.9M pixels) or in the number of generated tokens (SigLIP outputs an average of 2,410 tokens per sample, compared to 1,313 tokens for RADIO-AMP), we observe slightly improved accuracy with RADIO-AMP despite producing a significantly smaller number of output tokens.
Appendix A.2 Semantic Segmentation
Model Params ADE20k Pascal VOC
512 768 1024 512 768 1024
RADIO-AMP-B (ours) 98M 48.94 50.48 51.16 84.35 85.47 85.33
AM-RADIO-L [30]* 320M 50.03 37.99 35.63 83.76 68.85 63.86
+ multi-res 320M 51.54 51.74 52.84 85.51 86.84 87.21
RADIO-AMP-L (ours) 320M 51.47 51.90 52.95 85.49 86.96 87.03
AM-RADIO-H [30]* 553M 51.34 35.78 32.99 84.71 64.54 59.15
RADIO-AMP-H (ours) 553M 51.58 52.45 53.91 85.97 87.54 87.69
DINOv2-g-reg 1.1B 48.79 48.37 50.71 82.72 83.95 84.25
Table A5: Semantic segmentation mIoU for ADE20k and Pascal VOC across different models and resolutions: $512\times 512$, $768\times 768$, and $1024\times 1024$. Note: For DINOv2, we use the nearest larger multiple of its patch size (14). We apply a linear probe on top of the frozen features from the vision encoder. The mIoU of RADIO-AMP-B exceeds that of DINOv2-g-reg, despite being only one-tenth the size. *Our reproduction.
In Table A5, we report Semantic Segmentation mIoU at different resolutions on ADE20k [45] and VOC [12]. We observe that RADIO-AMP favorably scales to higher resolutions, while the accuracy of AM-RADIO falls above a resolution of $512\times 512$.
Appendix A.3 Additional Benchmarks
Following MLoRE [42] we report metrics on NYUDv2 and PASCAL Context for our model, along with the other agglomerative models (Theia [33] and UNIC [32]), as well as DINOv2 [9] as it’s the core perception teacher for these types of tasks. Previously to this study, MLoRE was the state-of-the-art for producing a multi-task model for these metrics. We keep the backbone model frozen, and use MLoRE’s “conv” head for each task. The conv head is defined as follows:
[…]
At the largest scale, RADIO-AMP-H is extremely competitive with DINOv2-g at half the parameters. At the ViT-B and ViT-L scale, RADIO-AMP is the closest to DINOv2 of the same size of any of the current agglomerative models. We train with a learning rate of $31e-$ 3, and use a weight of $1$ for all tasks. We purposefully don’t tune any hyperparameters.
Model Backbone SemSeg mIoU $\uparrow$ Depth RMSE $\downarrow$ Normal mErr $\downarrow$ Boundary Loss $\downarrow$
MLoRE Custom 55.96 0.5076 18.33 -
DINOv2 ViT-B/16 60.64 0.4816 18.19 0.1268
Theia ViT-B/16 38.90 0.6377 24.11 0.1298
UNIC ViT-B/16 42.21 0.6172 22.78 0.1285
RADIO-AMP ViT-B/16 57.19 0.4980 20.04 0.1263
DINOv2 ViT-L/14 62.94 0.4406 17.63 0.1266
UNIC ViT-L/14 58.56 0.4916 19.34 0.1274
RADIO-AMP ViT-L/16 61.42 0.4577 18.57 0.1259
AM-RADIO ViT-H/16 62.76 0.4339 18.43 0.1266
RADIO-AMP ViT-H/16 63.82 0.4353 17.67 0.1256
DINOv2 ViT-g/14 63.89 0.4252 17.20 0.1262
Table A6: Multi-task dense results on the NYUDv2 dataset. Note that we report the “Boundary Loss” and not the osdF metric due to the latter having a dependency on Matlab to compute.
Model Backbone SemSeg mIoU $\uparrow$ Parsing mIoU $\uparrow$ Saliency maxF $\uparrow$ Normal mErr $\downarrow$ Boundary Loss $\downarrow$
MLoRE Custom 81.41 70.52 84.90 13.51 -
DINOv2 ViT-B/16 81.68 73.24 77.54 17.44 0.0619
Theia ViT-B/16 69.84 60.67 80.63 16.94 0.0623
UNIC ViT-B/16 75.90 62.85 81.84 15.78 0.0620
RADIO-AMP ViT-B/16 81.75 71.49 81.26 16.10 0.0618
DINOv2 ViT-L/14 81.97 74.51 77.22 17.77 0.0620
UNIC ViT-L/14 81.82 72.24 79.21 17.35 0.0621
RADIO-AMP ViT-L/16 82.87 74.32 81.65 16.15 0.0617
AM-RADIO ViT-H/16 82.78 74.42 78.48 17.53 0.0619
RADIO-AMP ViT-H/16 83.43 75.75 81.19 16.16 0.0617
DINOv2 ViT-g/14 82.47 75.56 76.93 17.59 0.0619
Table A7: Multi-task dense results on the PASCAL Context dataset. Note that we report the “Boundary Loss” and not the osdF metric due to the latter having a dependency on Matlab to compute. We also don’t bold a cell in a model size group if a smaller model has even higher quality for a given benchmark.
Appendix A.4 Hyperparameters
Parameter Value
RADIO-AMP-L RADIO-AMP-H
Backbone ViT-L ViT-H
Learning Rate $1e^{-3}$
Weight Decay $2e^{-2}$
Teachers DFN CLIP
SigLIP 400m
DINOv2-g-reg
SAM-H
Feature Normalization PHI-S
Dataset DataComp-1B
Feature Distillation Loss MSE
Summary Loss Cosine
Backbone pre-training ImageNet-1k
Table A8: RADIO-AMP Training Hyperparameters
In Table A8, we report our RADIO-AMP training parameters.
Appendix A.5 A Measure of Scale Equivariance
We define a measure of scale equivariance $\sigma^{2}_{\text{scale}}$: given an array of feature tensors $\{F_{i}\}$ with shapes $(H_{i},W_{i},C)$:
1. Let $F_{\text{min}}$ denote the tensor with the smallest spatial dimensions $(H_{\text{min}},W_{\text{min}},C)$.
2. Compute the per-channel mean and variance of $F_{\text{min}}$:
$\mu_{c}=\frac{1}{H_{\text{min}}W_{\text{min}}}\sum_{h=1}^{H_{\text{min}}}\sum_{w=1}^{W_{\text{min}}}F_{\text{min}}(h,w,c)$ (4)
$\sigma_{c}^{2}=\frac{1}{H_{\text{min}}W_{\text{min}}}\sum_{h=1}^{H_{\text{min}}}\sum_{w=1}^{W_{\text{min}}}\left(F_{\text{min}}(h,w,c)-\mu_{c}\right)^{2}$ (5)
3. Normalize each tensor $F_{i}$ using $\mu_{c}$ and $\sigma_{c}$:
$\hat{F}_{i}(h,w,c)=\frac{F_{i}(h,w,c)-\mu_{c}}{\sigma_{c}}$ (6)
4. Bilinearly interpolate each normalized tensor $\hat{F}_{i}$ down to $(H_{\text{min}},W_{\text{min}},C)$, resulting in tensors $\{\tilde{F}_{i}\}$.
5. Stack all resized tensors $\{\tilde{F}_{i}\}$ along a new dimension and compute variance along this new dimension:
$\sigma^{2}(h,w,c)=\text{Var}(\{\tilde{F}_{i}(h,w,c)\})$ (7)
6. Finally, compute the average variance over the spatial dimensions:
$\sigma^{2}_{\text{scale}}=\frac{1}{H_{\text{min}}W_{\text{min}}}\sum_{h=1}^{H_{\text{min}}}\sum_{w=1}^{W_{\text{min}}}\sigma^{2}(h,w,c)$ (8)
A.5.1 Scale Variance Implementation
1
2def scale_variance(tensors: List, scale_up: bool):
3 """Compute feature variance across scales.
4
5 Steps:
6 * Per-channel standardization using the stats (mean/std)
7 of largest features (if scale_up) or smallest features (if scale_down)
8 * Interpolation to the size of the largest features (if scale_up) or
9 smallest features (if scale_down).
10 * Stack along a new dimension.
11 * Compute variance along the new dimension.
12 * Average across batch and spatial dimensions.
13 """
14
15 if scale_up:
16 target_tensor = max(tensors, key=lambda x: x.numel())
17 # Find the largest spatial dimensions
18 target_H = max(tensor.shape[1] for tensor in tensors)
19 target_W = max(tensor.shape[2] for tensor in tensors)
20 else:
21 target_tensor = min(tensors, key=lambda x: x.numel())
22 # Find the smallest spatial dimensions
23 target_H = min(tensor.shape[1] for tensor in tensors)
24 target_W = min(tensor.shape[2] for tensor in tensors)
25
26 # Compute mean and std along spatial dimensions (H, W) for each channel
27 mean = target_tensor.mean(dim=(1, 2), keepdim=True)
28 std = target_tensor.std(dim=(1, 2), keepdim=True)
29
30 # Normalize each tensor and resize to the largest spatial dimensions
31 normalized_tensors = []
32 for tensor in tensors:
33 # Mean-center and normalize
34 normalized_tensor = (tensor - mean) / (std + 1e-8) # Adding a small epsilon to avoid division by zero
35
36 # Resize to (B, max_H, max_W, C)
37 resized_tensor = F.interpolate(
38 normalized_tensor.permute(0, 3, 1, 2),
39 size=(target_H, target_W),
40 mode=’bilinear’,
41 align_corners=False
42 ).permute(0, 2, 3, 1)
43
44 normalized_tensors.append(resized_tensor)
45
46 stacked_tensors = torch.stack(normalized_tensors) # Shape: (num_tensors, B, max_H, max_W, C)
47
48 # Compute variance along the first dimension (num_tensors)
49 variance_tensor = torch.var(stacked_tensors, dim=0)
50
51 return variance_tensor.mean().item()
Appendix A.6 Mode-Switch PCA Visualizations
Figure A1: Visualizations of model features exhibiting the mode switch issue. We use PCA to project patch tokens into a 3D-space representing RGB colors. From left to right: input image, DINOv2, RADIO (baseline model) at 256x256, 384x384, 768x768, 1024x1024, and SAM. The visualizations illustrate how our baseline RADIO switches from producing DINO-like features at low resolution to producing SAM-like features at high resolution.
Figure A1 shows more visualization of the RADIO feature changes incurred by resolution increases.
Appendix A.7 Visualizations of Native vs Emulated High-Resolution Inference
Figure A2: Visualization of image upscaling with tiling and SigLIP
Figure A2 shows visualizations of output features for emulated high-resolution inference through tiling, and for native high-resolution inference using RADIO-AMP.
Appendix A.8 Token Merging
A.8.1 Ablation Study on ToMe Parameters
Sink Layout SigLIP RADIO-AMP-H
Values Keys Values
$4\times 4$ 0.56 0.54 0.50
$6\times 6$ 0.52 0.55 0.48
$8\times 8$ 0.56 0.59 0.53
Table A9: Reconstruction error (normalized MSE) after token merging/unmerging, using as criteria the "keys" (we use the attention keys) or "values" (we use the patch token values). The “Sink Layout” column indicates the strided arrangements for the merge sinks. RADIO-AMP-H exhibits a systematically lower error than SigLIP.
A.8.2 More Token Merging Visualizations
Figure A3: ToMe visualizations
Figure A3 shows more visualization of the ToME compression/decompression. Each input image yields 27x27=729 tokens, which are then compressed to 9 tokens using ToME.
Appendix A.9 Mosaic Visualizations
Figures A4 and A5 show visualizations of mosaic augmentations under $2\times 2$ and $4\times 4$ arrangements.
Figure A4: Mosaic visualization in a $4\times 4$ arrangement
Figure A5: Mosaic visualization in a $2\times 2$ arrangement